OpenStudy (anonymous):

Find cos(x/2) when sin(x)=-12/13, and x is in Quad III

OpenStudy (anonymous):

make the coreesponding triangle to find cos(x) then use the half angle identity.

OpenStudy (anonymous):



OpenStudy (anonymous):

a=-6.93?

OpenStudy (anonymous):

 \(a =-\sqrt{13^2-11^2}\)

OpenStudy (anonymous):

\(a = \sqrt{48}=\sqrt{16}\cdot\sqrt{3}=4\sqrt{3}\)

OpenStudy (anonymous):

oops... \(a=-4\sqrt{3}\)